A total of $10,000 is invested in two funds paying 5% and 7% annual interest. The combined interest is $630. How much of the $10,000 is invested in each fund?
Answer by checkley75(3666)

.07x+.05(10000-x)=630
.07x+500-.05x=630
.02x=630-500
.02x=130
x=130/.02
x=6500 invested @7%.
thus the remaining 10000-6500=3500 invested @5%.
proof
.07*6500+.05*3500=630
455+175=630
630=630